### v1

@@ 3.1.2 Estimates for offspring moments
In this part, we provide estimates for the offspring moments appearing in Corollary 3.3. We start with the $(2,2)$-type moment appearing in Corollary 3.3 (2). Recall the conditions (H1) and (H2”) from Section 1.1.
@@
For any $n$ sufficiently large, $\delta>0$ small and $\delta n\leq s\leq h(q_{n})-\delta n$,
${\mathbb{E}}_{n}[(\nu^{s}_{1})^{2}(\nu^{s}_{2})^{2}]\leq\frac{q_{n}(s+1)}{q_{n}(s)-1}{\mathbb{E}}_{n}[(\nu^{s}_{1})^{3}]$ (15)
In particular, under conditions (H1) and (H2”), for any $\delta>0$,
$\lim_{n\to\infty}\sup_{\delta n\leq s\leq h(q_{n})-\delta n}\frac{\mathbb{E}_{n}[(\nu_{1}^{s})^{2}(\nu_{2}^{s})^{2}]}{n}=0.$ (16)
Proof 3.6.
We observe that $(\nu^{s}_{1})^{2}(\nu^{s}_{2})^{2}\leq\nu^{s}_{1}(\nu^{s}_{2})^{3}+(\nu^{s}_{1})^{3}\nu^{s}_{2}$ and hence ${\mathbb{E}}_{n}[(\nu^{s}_{1})^{2}(\nu^{s}_{2})^{2}]\leq{\mathbb{E}}_{n}[(\nu^{s}_{1})^{3}\nu^{s}_{2}]$ by the exchangeability of $\nu$. In the same reason, we derive that
${\mathbb{E}}_{n}[(\nu^{s}_{1})^{3}\nu^{s}_{2}]=\frac{1}{q_{n}(s)-1}{\mathbb{E}}_{n}\Big[(\nu^{s}_{1})^{3}\sum_{j=2}^{q_{n}(s)}\nu^{s}_{j}\Big]$
$=\,$ $\frac{q_{n}(s+1)}{q_{n}(s)-1}{\mathbb{E}}_{n}[(\nu^{s}_{1})^{3}]-\frac{1}{q_{n}(s)-1}{\mathbb{E}}_{n}[(\nu^{s}_{1})^{4}]\leq\frac{q_{n}(s+1)}{q_{n}(s)-1}{\mathbb{E}}_{n}[(\nu^{s}_{1})^{3}].$
Therefore (15) follows. Under assumption (H1), since $\ell$ is continuous, it is easy to obtain that for all $\delta>0$, $\sup_{\delta n\leq s\leq h(q_{n})-\delta n}\frac{q_{n}(s+1)}{q_{n}(s)-1}\to 1$ as $n\to\infty$. (16) is then a consequence of (15) and (H2”).
Next, we consider the moments in Corollary 3.3 (1) with $l=1,2$.
@@
Assume (H1) and (H2”). For any $\delta>0$ and $m\geq 1$, uniformly for $\delta n\leq s\leq h(q_{n})-\delta n$, we have:
  • (1)
    $\mathbb{E}_{n}\Big[\prod_{i=1}^{m}\nu_{i}^{s}\Big]=1+o_{n}(1),$ (17)
@@
    $\mathbb{E}_{n}\Big[\prod_{i=1}^{m}\nu_{i}^{s}\Big]=\frac{(q_{n}(s+1))^{m}}{(q_{n}(s))_{m}}-\dfrac{\binom{m}{2}\big(1+\sigma_{n}^{2}(s)\big)}{q_{n}(s)}\big(1+o_{n}(1)\big);$ (18)
  • (2)
    $\mathbb{E}_{n}\Big[(\nu_{1}^{s})^{2}\prod_{i=2}^{m}\nu_{i}^{s}\Big]=1+\sigma_{n}(s)^{2}+o_{n}(1).$ (19)
Proof 3.8.
@@
By conditioning on $\nu_{1}^{s},\dots,\nu_{m-1}^{s}$, we obtain by induction that
$a_{m}$ $=\mathbb{E}_{n}\Bigg[\frac{q_{n}(s+1)-\sum_{i=1}^{m-1}\nu_{i}^{s}}{q_{n}(s)-(m-1)}\prod_{i=1}^{m-1}\nu_{i}^{s}\Bigg]=\frac{q_{n}(s+1)}{q_{n}(s)-(m-1)}a_{m-1}-\frac{m-1}{q_{n}(s)-(m-1)}b_{m-1}$
$=\cdots=\frac{(q_{n}(s+1))^{m}}{(q_{n}(s))_{m}}-\sum_{i=1}^{m-1}\frac{(q_{n}(s+1))^{m-i-1}}{(q_{n}(s)-i)_{m-i}}\cdot ib_{i},\quad m\geq 1.$ (20)
@@
$b_{m}\leq\frac{(q_{n}(s+1))^{m-1}}{(q_{n}(s)-1)_{m-1}}b_{1}.$ (22)
Combining (3.8) and (22) gives that
@@
where both $o_{n}(1)$ and $O_{n}(\tfrac{1}{n})$ are uniform for $\delta n\leq s\leq h(q_{n})-\delta n$. Hence, (17) follows.
The bound (19) follows similarly. Indeed, in the same way as (22),
@@
Approximating $b_{1}$ by (H1), $c_{1}$ by (H2”) and $d_{2}$ by Lemma 3.5, we deduce (19) from (21). Finally, by (H1), uniformly over $i=1,\cdots,m-1$,
@@
and thus (18) follows directly from (19) together with (3.8).
@@
Assume (H1) and (H2”). For any $\delta>0$ and $m\geq 1$, uniformly for $\delta n\leq s\leq h(q_{n})-\delta n$, we have:
  • (1)
    ${\mathbb{P}}_{n}\Big(\mathfrak{p}(u_{1}^{s+1}),\ldots,\mathfrak{p}(u_{m}^{s+1})~\text{are distinct}\Big)=1+o_{n}(1),$
    and more precisely,
    $\begin{aligned} {\mathbb{P}}_{n}\Big(\mathfrak{p}(u_{1}^{s+1}),\ldots,\mathfrak{p}(u_{m}^{s+1})~\text{are distinct}\Big)=1-\dfrac{\binom{m}{2}\sigma_{n}(s)^{2}}{q_{n}(s)}\big(1+o_{n}(1)\big);\end{aligned}$ (23)
  • (2)
    $\begin{aligned} &{\mathbb{P}}_{n}\Big(\{\mathfrak{p}(u_{1}^{s+1})=\mathfrak{p}(u_{2}^{s+1})\}\cap\{\mathfrak{p}(u_{2}^{s+1}),\ldots,\mathfrak{p}(u_{m}^{s+1})~\text{are distinct}\}\Big)\\ =\,&\frac{1}{q_{n}(s)}\big(\sigma_{n}(s)^{2}+o_{n}(1)\big).\end{aligned}$ (24)
Proof 3.10.
Recall that by (H1), $\frac{q_{n}(s+1)}{q_{n}(s)}=1+o(1)$ uniformly for $\delta n\leq s\leq h(q_{n})-\delta n$. Hence by Corollary 3.3 (with $l=1$) and (18), we deduce that
@@
which yields (23).
Moreover,
@@
Therefore (24) follows from Corollary 3.3 (with $l=2$) and (19).
@@ 3.2 Coalescent process
In this subsection, we define the coalescent process and derive estimates for its transition probabilities.
We begin with the relevant notation. For any vertex $v\in\mathscr{T}^{n}$ and $0\leq i\leq|v|$, let $\mathrm{Anc}(v,i)$ denote the ancestor of $v$ at height $i$. Define the set of descendants of $v$ by
@@
Fix a height $0\leq h^{*}\leq h(q_{n})$ and $1\leq k\leq q_{n}(h^{*})$. For $0\leq j\leq h^{*}$, set
$\mathcal{X}^{(k)}_{j}=\mathcal{X}^{(k)}_{j}(n,h^{*})$ $=\big\{\text{Anc}(u^{h^{*}}_{i},j):1\leq i\leq k\big\}$ (25)
@@
In other words, $X^{(k)}_{j}$ represents the number of distinct ancestral lineages at height $j$ traced from the $k$ uniformly sampled vertices at height $h^{*}$. At $j=h^{*}$, we have $X^{(k)}_{0}=k$, since all sampled vertices are distinct. As $j$ decreases, the lineages are traced backward through the tree and may coalesce whenever two or more sampled vertices share a common ancestor. Thus, the process $\big(X^{(k)}_{j}:0\leq j\leq h^{*}\big)$ is non-increasing as $j$ decreases, and each downward jump corresponds to a coalescence event between ancestral lineages.
For each $v\in\mathscr{T}^{n}$ with $|v|\leq h^{*}$, we write
$\mathscr{T}^{n}(v):=\mathscr{T}^{n}(v,h^{*})$ (26)
for the subtree of $\mathscr{T}^{n}$ rooted at $v$, truncated at height $h^{*}$; that is, the subtree consisting of vertices
$\{u\in\operatorname{Desc}(v):|u|\leq h^{*}\}.$
We then define, for $0\leq j\leq h^{*}$ and $1\leq k\leq q_{n}(h^{*})$,
$\mathcal{G}^{(k)}_{j}=\mathcal{G}^{(k)}_{j}(h^{*}):=\sigma\!\big(\{\mathscr{T}^{n}(v):|v|=j\}\big)\vee\sigma(\{u^{h^{*}}_{i}:1\leq i\leq k\}).$ (27)
We emphasize that both collections $\{\mathscr{T}^{n}(v):|v|=j\}$ and $\{u^{h^{*}}_{i}:1\leq i\leq k\}$ are treated as unordered sets; in particular, $\mathcal{G}^{(k)}_{j}$ does not record any ordering among vertices at height $j$. When $k=q_{n}(h^{*})$, we write simply $\mathcal{G}_{j}=\mathcal{G}^{(k)}_{j}$, in which case
$\mathcal{G}_{j}=\sigma\!\big(\{\mathscr{T}^{n}(v):|v|=j\}\big).$
@@
Proof 3.13.
By the exchangeability of the offspring vector $\nu^{j}=\big(\nu_{1}^{j},\ldots,\nu_{q_{n}(j)}^{j}\big)$, the vector of subtrees
@@
is also exchangeable, a fact that can be readily verified. The conclusion then follows immediately.
@@
${\mathbb{P}}_{n}\big(X^{(k)}_{j-1}=m_{2}\,\big|\,X^{(k)}_{j}=m_{1}\big)={\mathbb{P}}_{n}\!\left(\#\{\mathfrak{p}(u^{j}_{i}):1\leq i\leq m_{1}\}=m_{2}\right).$ (28)
@@
$\mathbb{E}\big(f(X,Y)\mid Z\big)=\mathbb{E}\big(f(W,Y)\big).$ (30)
Proof 3.16.
By the standard property of conditional expectation for independent random elements (see, e.g., [durrett2019probability, Example 4.1.7]),
@@
Applying the tower property of conditional expectation yields (29).
For (30), note that
@@
Let $1\leq m_{2}\leq m_{1}\leq q_{n}(j)$ and $k_{1},\ldots,k_{m_{2}}\in[1,q_{n}(j-1)]$ be distinct. Then on $\{X_{j}=m_{1}\}$,
$\mathbb{P}_{n}\big(\mathcal{X}_{j}=\{x^{j-1}_{k_{1}},\ldots,x^{j-1}_{k_{m_{2}}}\}\,\big|\,\mathcal{G}_{j}\big)=\binom{q_{n}(j-1)}{m_{2}}^{-1}{\mathbb{P}}_{n}\big(\#\{\mathfrak{p}(u^{j-1}_{i}):1\leq i\leq m_{1}\}=m_{2}\big).$ (31)
Proof 3.18.
@@
$\begin{aligned} &\mathbb{P}_{n}\big(\mathcal{X}_{j}=\{x^{j-1}_{k_{1}},\ldots,x^{j-1}_{k_{m_{2}}}\}\,\big|\,\mathcal{G}_{j}\big)\\ =\;&{\mathbb{P}}_{n}\Big(\big\{\mathfrak{p}(x):x\in\mathcal{X}_{j}\big\}=\{x^{j-1}_{k_{1}},\ldots,x^{j-1}_{k_{m_{2}}}\}\,\big|\,\big(\mathscr{T}^{n}(v):|v|=j\big)\Big).\end{aligned}$ (32)
@@
    Both $\mathcal{X}_{j}$ and $\big(\mathscr{T}^{n}(v):|v|=j\big)$ are measurable with respect to
    $\sigma\big(\nu^{j},\nu^{j+1}\ldots,\nu^{h^{*}}\big),$
    and are independent of $\nu^{j-1}$ due to generation-wise independence of $\nu^{s}$.
@@
    The event $\big\{\mathfrak{p}(x):x\in\mathcal{X}_{j}\big\}=\{x^{j-1}_{k_{1}},\ldots,x^{j-1}_{k_{m_{2}}}\}$ is determined by $\mathcal{X}_{j}$ and $\nu^{j-1}$, i.e.,
    $\Big\{\big\{\mathfrak{p}(x):x\in\mathcal{X}_{j}\big\}=\{x^{j-1}_{k_{1}},\ldots,x^{j-1}_{k_{m_{2}}}\}\Big\}\in\sigma(\mathcal{X}_{j},\nu^{j-1}).$
@@
$X=\mathcal{X}_{j},\quad Y=\nu^{j-1},\quad Z=\big(\mathscr{T}^{n}(v):|v|=j\big),\quad W=\{u^{j}_{1},\ldots,u^{j}_{m_{1}}\},$
and letting $f(X,Y)$ be the function determined by $\mathcal{X}_{j-1}$ and $\nu^{d_{j}}$ corresponding to the event
@@
we obtain that on $\{X_{j}=m_{1}\}$,
${\mathbb{P}}_{n}\Big(\big\{\mathfrak{p}(x):x\in\mathcal{X}_{j}\big\}=\{x^{j-1}_{k_{1}},\ldots,x^{j-1}_{k_{m_{2}}}\}\,\big|\,\big(\mathscr{T}^{n}(v):|v|=j\big)\Big)$ (33)
@@
Substituting back into (32) gives the desired conclusion.
Proof 3.19 (Proof of Theorem 3.14).
It follows from (31) that on $\{X_{j}=m_{1}\}$,
${\mathbb{P}}_{n}\big(X_{j-1}=m_{2}\,\big|\,\mathcal{G}_{j}\big)$ $=\sum_{1\leq k_{1}<\cdots<k_{m_{1}}\leq q_{n}(j-1)}\mathbb{P}_{n}\big(\mathcal{X}_{j}=\{x^{j-1}_{k_{1}},\ldots,x^{j-1}_{k_{m_{1}}}\}\,\big|\,\mathcal{G}_{j}\big)$
@@
This confirms the Markov property and the transition formula (28).
We close this section by a simple estimate.
@@
${\mathbb{P}}_{n}\big(X^{(k)}_{j-1}\leq m-1\,\big|\,X^{(k)}_{j}=m\big)\lesssim\frac{m^{2}}{n}.$ (34)
Proof 3.21.
By Corollary 3.3 and Theorem 3.14,

### v2

@@ 3.1.2 Estimates for offspring moments
In this part, we provide estimates for the offspring moments appearing in Corollary 3.2. We start with the $(2,2)$-type moment appearing in (3.5). Recall the conditions (H1) and (H2b) from Section 1.1.
@@
For any $n\geq 1$ and $1\leq s<h(q_{n})$,
${\mathbb{E}}_{n}[(\nu^{s}_{1})^{2}(\nu^{s}_{2})^{2}]\leq\frac{q_{n}(s+1)}{q_{n}(s)-1}{\mathbb{E}}_{n}[(\nu^{s}_{1})^{3}]$ (3.6)
In particular, under conditions (H1) and (H2b), for any $\delta>0$,
$\lim_{n\to\infty}\sup_{\delta n\leq s\leq h(q_{n})-\delta n}\frac{\mathbb{E}_{n}[(\nu_{1}^{s})^{2}(\nu_{2}^{s})^{2}]}{n}=0.$ (3.7)
Proof.
We first note the elementary inequality $2(\nu^{s}_{1})^{2}(\nu^{s}_{2})^{2}\leq\nu^{s}_{1}(\nu^{s}_{2})^{3}+(\nu^{s}_{1})^{3}\nu^{s}_{2}$. By exchangeability of the offspring vector $\nu^{s}$, this implies ${\mathbb{E}}_{n}[(\nu^{s}_{1})^{2}(\nu^{s}_{2})^{2}]\leq{\mathbb{E}}_{n}[(\nu^{s}_{1})^{3}\nu^{s}_{2}]$. Moreover,
${\mathbb{E}}_{n}[(\nu^{s}_{1})^{3}\nu^{s}_{2}]=\frac{1}{q_{n}(s)-1}{\mathbb{E}}_{n}\Big[(\nu^{s}_{1})^{3}\sum_{j=2}^{q_{n}(s)}\nu^{s}_{j}\Big]=\frac{q_{n}(s+1)}{q_{n}(s)-1}{\mathbb{E}}_{n}[(\nu^{s}_{1})^{3}]-\frac{1}{q_{n}(s)-1}{\mathbb{E}}_{n}[(\nu^{s}_{1})^{4}]\leq\frac{q_{n}(s+1)}{q_{n}(s)-1}{\mathbb{E}}_{n}[(\nu^{s}_{1})^{3}],$
which proves (3.6).
Under condition (H1), the continuity of $\ell$ implies that for every $\delta>0$,
$\sup_{\delta n\leq s\leq h(q_{n})-\delta n}\frac{q_{n}(s+1)}{q_{n}(s)-1}\;\longrightarrow\;1\qquad\text{as }n\to\infty.$
Combining this with (3.6) and (H2b) yields (3.7). ∎
Next, we consider the moments in (3.4) with $l=1,2$.
@@
Assume (H1) and (H2b). For any $\delta>0$ and $m\geq 1$, uniformly for $\delta n\leq s\leq h(q_{n})-\delta n$, we have
$\mathbb{E}_{n}\Big[\prod_{i=1}^{m}\nu_{i}^{s}\Big]=1+o_{n}(1),$ (3.8)
@@
$\mathbb{E}_{n}\Big[\prod_{i=1}^{m}\nu_{i}^{s}\Big]=\frac{(q_{n}(s+1))^{m}}{(q_{n}(s))_{m}}-\dfrac{\binom{m}{2}\big(1+\sigma_{n}^{2}(s)\big)}{q_{n}(s)}\big(1+o_{n}(1)\big);$ (3.9)
Furthermore,
$\mathbb{E}_{n}\Big[(\nu_{1}^{s})^{2}\prod_{i=2}^{m}\nu_{i}^{s}\Big]=1+\sigma_{n}(s)^{2}+o_{n}(1).$ (3.10)
Proof.
@@
By conditioning on $\nu_{1}^{s},\dots,\nu_{m-1}^{s}$, we obtain by induction that for any $m\geq 1$
$\begin{aligned} a_{m}&=\mathbb{E}_{n}\Bigg[\frac{q_{n}(s+1)-\sum_{i=1}^{m-1}\nu_{i}^{s}}{q_{n}(s)-(m-1)}\prod_{i=1}^{m-1}\nu_{i}^{s}\Bigg]=\frac{q_{n}(s+1)}{q_{n}(s)-(m-1)}a_{m-1}-\frac{m-1}{q_{n}(s)-(m-1)}b_{m-1}\\ &=\cdots=\frac{(q_{n}(s+1))^{m}}{(q_{n}(s))_{m}}-\sum_{i=1}^{m-1}\frac{(q_{n}(s+1))^{m-i-1}}{(q_{n}(s)-i)_{m-i}}\cdot ib_{i},\end{aligned}$ (3.11)
@@
$b_{m}\leq\frac{(q_{n}(s+1))^{m-1}}{(q_{n}(s)-1)_{m-1}}b_{1}.$ (3.13)
Combining (3.11) and (3.13), we derive that
@@
where both $o_{n}(1)$ and $O_{n}(\tfrac{1}{n})$ are uniform in $\delta n\leq s\leq h(q_{n})-\delta n$. Hence, (3.8) follows.
The bound (3.10) follows similarly. Indeed, in the same way as (3.13),
@@
Approximating $b_{1}$ by (H1), $c_{1}$ by (H2b) and $d_{2}$ by Lemma 3.4, we deduce (3.10) from (3.12). Finally, by (H1), uniformly over $i=1,\cdots,m-1$,
@@
and thus (3.9) follows directly from (3.10) together with (3.11). ∎
@@
Assume (H1) and (H2b). For any $\delta>0$ and $m\geq 1$, uniformly for $\delta n\leq s\leq h(q_{n})-\delta n$, we have:
${\mathbb{P}}_{n}\Big(\mathfrak{p}(u_{1}^{s+1}),\ldots,\mathfrak{p}(u_{m}^{s+1})~\text{are distinct}\Big)=1-\dfrac{\binom{m}{2}\sigma_{n}(s)^{2}}{q_{n}(s)}\big(1+o_{n}(1)\big);$ (3.14)
${\mathbb{P}}_{n}\Big(\{\mathfrak{p}(u_{1}^{s+1})=\mathfrak{p}(u_{2}^{s+1})\}\cap\{\mathfrak{p}(u_{2}^{s+1}),\ldots,\mathfrak{p}(u_{m}^{s+1})~\text{are distinct}\}\Big)=\,\frac{1}{q_{n}(s)}\big(\sigma_{n}(s)^{2}+o_{n}(1)\big).$ (3.15)
Proof.
Recall that by (H1), $\frac{q_{n}(s+1)}{q_{n}(s)}=1+o_{n}(1)$ uniformly over $\delta n\leq s\leq h(q_{n})-\delta n$. By (3.4) (with $l=1$) in Corollary 3.2 and (3.9), we deduce that
@@
which yields (3.14). Moreover, by (3.8) and (3.10),
@@
Therefore (3.15) follows from (3.4) (with $l=2$). ∎
@@ 3.2 Coalescent process
In this subsection, we introduce the coalescent process and study its transition probabilities. The central result is Theorem 3.9, which identifies the coalescent process as a (time-inhomogeneous) Markov chain and provides its transition formula.
We begin with some relevant notation. For any vertex $v\in\mathscr{T}^{n}$ and $0\leq i\leq|v|$, let $\mathrm{Anc}(v,i)$ denote the ancestor of $v$ at height $i$. Define the set of descendants of $v$ by
@@
Fix a height $0\leq h^{*}\leq h(q_{n})$ and an integer $1\leq k\leq q_{n}(h^{*})$. For $0\leq j\leq h^{*}$, set
$\mathcal{X}^{(k)}_{j}=\mathcal{X}^{(k)}_{j}(n,h^{*})$ $=\big\{\text{Anc}(u^{h^{*}}_{i},j):1\leq i\leq k\big\}$ (3.16)
@@
In other words, $X^{(k)}_{j}$ represents the number of distinct ancestral lineages at height $j$ traced from $k$ uniformly sampled vertices at height $h^{*}$. At $j=h^{*}$, we have $X^{(k)}_{0}=k$, since all sampled vertices are distinct. As $j$ decreases, the lineages are traced backward through the tree and may coalesce whenever two or more sampled vertices share a common ancestor. Thus, the process $\big(X^{(k)}_{j}:0\leq j\leq h^{*}\big)$ is non-increasing as $j$ decreases, and each downward jump corresponds to a coalescence event between ancestral lineages.
We consider subtrees of $\mathscr{T}^{n}$ rooted at some vertex $v$ and truncated at height $h^{*}$. To make this precise, we recall the representation of a tree as a subset of $\mathcal{U}$ in Section 2.1. For $u=(u_{1},\ldots,u_{m})$, $v=(v_{1},\ldots,v_{l})\in\mathcal{U}$, we write $uv$ for the concatenated sequence $(u_{1},\ldots,u_{m},v_{1},\ldots,v_{l})$. For each $v\in V(\mathscr{T}^{n})$, define the truncated subtree
$\mathscr{T}^{n}(v)=\mathscr{T}^{n}(v,h^{*}):=\big\{u\in\mathcal{U}:vu\in\mathscr{T}^{n},\;\left\lvert vu\right\rvert\leq h^{*}\big\}.$ (3.17)
Moreover, for any $1\leq k\leq q_{n}(h^{*})$ and $v\in V(\mathscr{T}^{n})$, we define the truncated subtree consisting only of the ancestral lineages of the sampled vertices $\{u^{h^{*}}_{i}:1\leq i\leq k\}$:
$\mathscr{T}^{n,(k)}(v)=\mathscr{T}^{n,(k)}(v,h^{*}):=\big\{u\in\mathcal{U}:vu\in\mathscr{T}^{n}(v),\;\exists\,1\leq i\leq k\text{ s.t. }u^{h^{*}}_{i}\in\text{Desc}(vu)\big\}.$
Note that both $\mathscr{T}^{n}(v)$ and $\mathscr{T}^{n,(k)}(v)$ record only the suffix $u$ in the decomposition $vu$; the prefix $v$ itself is forgotten in this representation. For $0\leq j\leq h^{*}$ and $1\leq k\leq q_{n}(h^{*})$, set
$\mathcal{G}^{(k)}_{j}=\mathcal{G}^{(k)}_{j}(h^{*}):=\sigma\!\big(\big\{\big(\mathscr{T}^{n}(v),\mathscr{T}^{n,(k)}(v)\big):\left\lvert v\right\rvert=j\big\}\big).$ (3.18)
Here the collection $\big\{\big(\mathscr{T}^{n}(v),\mathscr{T}^{n,(k)}(v)\big):|v|=j\big\}$ is regarded as a set (rather than a sequence), so no ordering information at generation $j$ is retained. Clearly, as $j$ decreases, the $\sigma$-fields $\mathcal{G}^{(k)}_{j}$ form an increasing family. See Figure 2 for an illustration of the $\sigma$-fields.
When $k=q_{n}(h^{*})$, we simply write $\mathcal{G}_{j}=\mathcal{G}^{(k)}_{j}$; in this case, since $\mathscr{T}^{n}(v)$ already determines $\mathscr{T}^{n,(q_{n}(h^{*}))}(v)$.
$\mathcal{G}_{j}=\sigma\!\big(\{\mathscr{T}^{n}(v):\left\lvert v\right\rvert=j\}\big),$ (3.19)
Figure 2: Illustration of the $\sigma$-field $\mathcal{G}^{(k)}_{j}$ with $h^{*}=3$, $j=1$ and $k=5$. The red points represent the $k=5$ uniformly sampled vertices at height $h^{*}=3$. For each $\left\lvert v\right\rvert=1$, the tree $\mathscr{T}^{n}(v)$ consists of the entire descendant subtree of $v$ up to height $h^{*}$ (shown in both the black and blue), while $\mathscr{T}^{n,(k)}(v)$ (blue only) records the portions of $\mathscr{T}^{n}(v)$ that lie on the ancestral lineages of the sampled vertices. The $\sigma$-field $\mathcal{G}^{(k)}_{j}$ is generated by these subtrees regarded as a set.
@@
Proof.
By the exchangeability of the offspring vector $\nu^{j}=\big(\nu_{1}^{j},\ldots,\nu_{q_{n}(j)}^{j}\big)$, we can readily verify that the law of the vector of subtrees
@@
is also exchangeable. The claimed conditional distribution of $\mathcal{X}^{(k)}_{j}$ now easily follows. ∎
@@
${\mathbb{P}}_{n}\big(X^{(k)}_{j-1}=m_{2}\,\big|\,X^{(k)}_{j}=m_{1}\big)={\mathbb{P}}_{n}\!\left(\#\{\mathfrak{p}(u^{j}_{i}):1\leq i\leq m_{1}\}=m_{2}\right).$ (3.20)
@@
$\mathbb{E}\big(f(X,Y)\mid Z\big)=\mathbb{E}\big(f(W,Y)\big).$ (3.22)
Proof.
By the standard property of conditional expectation for independent random elements (see, e.g., [7, Example 4.1.7]),
@@
Applying the tower property of conditional expectation yields (3.21). For (3.22), note that
@@
Let $1\leq m_{2}\leq m_{1}\leq q_{n}(j)$ and $k_{1},\ldots,k_{m_{2}}\in[1,q_{n}(j-1)]$ be distinct. Then on the event $\{X_{j}=m_{1}\}$,
$\mathbb{P}_{n}\big(\mathcal{X}_{j-1}=\{x^{j-1}_{k_{1}},\ldots,x^{j-1}_{k_{m_{2}}}\}\,\big|\,\mathcal{G}_{j}\big)=\binom{q_{n}(j-1)}{m_{2}}^{-1}{\mathbb{P}}_{n}\big(\#\{\mathfrak{p}(u^{j}_{i}):1\leq i\leq m_{1}\}=m_{2}\big).$ (3.23)
Proof.
@@
$\begin{aligned} \mathbb{P}_{n}\big(\mathcal{X}_{j-1}=\{x^{j-1}_{k_{1}},\ldots,x^{j-1}_{k_{m_{2}}}\}\,\big|\,\mathcal{G}_{j}\big)=\;{\mathbb{P}}_{n}\Big(\big\{\mathfrak{p}(x):x\in\mathcal{X}_{j}\big\}=\{x^{j-1}_{k_{1}},\ldots,x^{j-1}_{k_{m_{2}}}\}\,\big|\,\big(\mathscr{T}^{n}(v):|v|=j\big)\Big).\end{aligned}$ (3.24)
@@
    Both $\mathcal{X}_{j}$ and $\big(\mathscr{T}^{n}(v):|v|=j\big)$ are measurable with respect to $\sigma\big(\nu^{j},\nu^{j+1}\ldots,\nu^{h^{*}}\big),$ therefore independent of $\nu^{j-1}$ due to generation-wise independence of $\nu^{s}$.
@@
    The event $\Big\{\big\{\mathfrak{p}(x):x\in\mathcal{X}_{j}\big\}=\big\{x^{j-1}_{k_{1}},\ldots,x^{j-1}_{k_{m_{2}}}\big\}\Big\}$ is determined by $\mathcal{X}_{j}$ and $\nu^{j-1}$, i.e.,
    $\Big\{\big\{\mathfrak{p}(x):x\in\mathcal{X}_{j}\big\}=\big\{x^{j-1}_{k_{1}},\ldots,x^{j-1}_{k_{m_{2}}}\big\}\Big\}\in\sigma(\mathcal{X}_{j},\nu^{j-1}).$
@@
$X=\mathcal{X}_{j},\quad Y=\nu^{j-1},\quad Z=\big\{\mathscr{T}^{n}(v):|v|=j\big\},\quad W=\{u^{j}_{1},\ldots,u^{j}_{m_{1}}\},$
and $f(X,Y)$ the indicator function determined by $\mathcal{X}_{j}$ and $\nu^{j-1}$ corresponding to the event
@@
we obtain that on the event $\{X_{j}=m_{1}\}$,
${\mathbb{P}}_{n}\Big(\big\{\mathfrak{p}(x):x\in\mathcal{X}_{j}\big\}=\{x^{j-1}_{k_{1}},\ldots,x^{j-1}_{k_{m_{2}}}\}\,\big|\,\big(\mathscr{T}^{n}(v):|v|=j\big)\Big)$ (3.25)
@@
Substituting back into (3.24) gives the desired conclusion. ∎
Proof of Theorem 3.9.
It follows from (3.23) that on $\{X_{j}=m_{1}\}$,
${\mathbb{P}}_{n}\big(X_{j-1}=m_{2}\,\big|\,\mathcal{G}_{j}\big)$ $=\sum_{1\leq k_{1}<\cdots<k_{m_{1}}\leq q_{n}(j-1)}\mathbb{P}_{n}\big(\mathcal{X}_{j-1}=\{x^{j-1}_{k_{1}},\ldots,x^{j-1}_{k_{m_{1}}}\}\,\big|\,\mathcal{G}_{j}\big)$
@@
This confirms the Markov property and the transition formula (3.20). ∎
We close this section with a simple estimate.
@@
${\mathbb{P}}_{n}\big(X^{(k)}_{j-1}\leq m-1\,\big|\,X^{(k)}_{j}=m\big)\lesssim\frac{m^{2}}{n}.$ (3.26)
Proof.
By (3.4) in Corollary 3.2 and Theorem 3.9,